OpenStudy (vshiroky):

For the equation 8y=4+x answer part A and B The x intercept is ____ The y intercept is ____

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (vshiroky):

I'm unsure of how to do this

OpenStudy (vshiroky):

I think 8y=4+x so divide by 8? y=4/8+x?

OpenStudy (skullpatrol):

What are the definitions of the "intercepts" ?

OpenStudy (vshiroky):

where they intersect eachother

OpenStudy (skullpatrol):

It is where the graph intersects the axis.

OpenStudy (vshiroky):

(-4,0) and (0,1/2)?

OpenStudy (vshiroky):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Those are both correct. Do you see how those are the x and y intercepts?

OpenStudy (skullpatrol):

The x intercept is ____ the value of y when x = 0 8y=4+x 8y=4+0 y=1/2

OpenStudy (vshiroky):

I do understand now... Is it -1/2 or 1/2?

jimthompson5910 (jim_thompson5910):

as skullpatrol is showing, it's positive The graph is attached

OpenStudy (anonymous):

x - 8y = 4; turn the equation in the slopeis the y-intercept. x - 8y = 4 x - x - 8y = -x + 4 -8y = -x + 4 divide by -8-intercept form, y = mx + b, where bto both sides y = x/8 - 1/2 Thus, the y-intercept is -1/2.